### v1

# Attribute Structuring Improves LLM-Based Evaluation of Clinical Text Summaries

Abstract
Summarizing clinical text is crucial in health decision-support and clinical research. Large language models (LLMs) have shown the potential to generate accurate clinical text summaries, but still struggle with issues regarding grounding and evaluation, especially in safety-critical domains such as health. Holistically evaluating text summaries is challenging because they may contain unsubstantiated information. Here, we explore a general mitigation framework using Attribute Structuring (AS), which structures the summary evaluation process. It decomposes the evaluation process into a grounded procedure that uses an LLM for relatively simple structuring and scoring tasks, rather than the full task of holistic summary evaluation. Experiments show that AS consistently improves the correspondence between human annotations and automated metrics in clinical text summarization. Additionally, AS yields interpretations in the form of a short text span corresponding to each output, which enables efficient human auditing, paving the way towards trustworthy evaluation of clinical information in resource-constrained scenarios. We release our code, prompts, and an open-source benchmark at https://github.com/microsoft/attribute-structuring/.
[…]
(Hide=100, Hang=true, Progressive=2ex, Style*$\bullet$ , Style2*$\circ$)
Attribute Structuring Improves LLM-Based Evaluation
of Clinical Text Summaries
Zelalem Gero, Chandan Singh, Yiqing Xie, Sheng Zhang, Tristan Naumann, Jianfeng Gao, Hoifung Poon Microsoft Research
1 Introduction
Automatic medical text summarization has emerged as a critical task for real-world impact in natural-language processing, amidst widespread physician fatigue and burnout. Meanwhile, pre-trained large language models (LLMs) have shown impressive proficiency in a range of complex natural language processing tasks [1, 2, 3], including in text summarization [4, 5]. However, evaluating these summarizations is challenging, as they rely on human effort to track different attributes and facts.
Here, we propose using LLMs for evaluating medical text summaries, guided by a simple structuring process. LLMs have previously been used to provide practical evaluations in a variety of evaluation settings [6, 7]. However, these evaluations have been imperfect: yielding relatively low correlation with human judgments, showing bias [8, 9], or more generally struggling with long-form nuanced evaluation, as is often the case in summarization [10]. Structuring the scoring process can help to ground the evaluation, and help an LLM decompose a complex evaluation process into simpler, more accurate steps.
We introduce a simple structuring process, which we call Attribute Structuring (AS) (Fig. 1). Given a groundtruth summary and a summary to be evaluated, we leverage an LLM to score a summary through prompting. Rather than directly asking the LLM for a score, we ask it to provide a score for each pair of attributes, one from the groundtruth and another from the summary to be evaluated. We extract 17 attributes derived from a clinical ontology for each summary; the attributes are shown to define the elements of quality in the context of a clinical discharge summary in a large scale study [11]. Unlike previous methods such as FActScore [12], this does not require a reliable knowledge source to verify each attribute, but instead relies on the pre-specified ontology to capture relevant aspects of a summary. DocLens  [13], similar to our work, performs a fine-grained evaluation of medical text using various generation models and show that such scoring has higher agreement with the judgements of medical experts. However, DocLens relies on the model to generate the important attributes rather than a clinical ontology.
Figure 1: Overview of Attribute Structuring LLM evaluation for clinical summarization. AS uses a human-given ontology to extract attributes from summaries and then to individually score them.
We perform experiments on clinical summarization tasks using discharge notes from MIMIC-III [14] and various LLMs, including GPT-4 [3], GPT-3.5 [15], and LLaMA-2 [2]. We find that that Attribute Structuring helps reduce the gap between automated metrics and human annotators. In addition, AS enables auditing the evaluation for a particular summary, by providing grounding for each attribute, in the form of a short text span in the input.
2 Methods and experimental setup
2.1 Methods: Attribute Structuring
Fig. 1 shows the various steps of the Attribute Structuring (AS) pipeline. The pipeline takes in two summaries: a groundtruth summary and a generated summary (which we are trying to evaluate). Additionally, it uses a set of attributes that are important for the task at hand. These attributes comes from an ontology that utilizes domain knowledge to define the elements of quality, in our case the context of a clinical discharge summary in a large scale study [11], e.g. Admission diagnosis or Goals of care.
Each step in the pipeline is achieved by prompting an LLM. In the first step, the Attribute Structuring pipeline first extracts text for each attribute in the ontology from both summaries. Then, for each pair of extracted attributes, it generates a similarity score by prompting an LLM to compare them on a scale of 1–5 (all prompts are provided in Appendix A). Finally, an average of these scores is returned as the evaluation score for the generated summary (usually after normalizing to range from 0–100) .
In a final (optional) interpretation step, an LLM can be used to ground the scores for each attribute using text spans in the input. This can be achieved by directly prompting an LLM to find a text span that led to the extracted text for each attribute. This step is made possible by the Attribute Structuring structuring process.
2.2 Experimental setup
Dataset and benchmark
We construct a benchmark consisting of discharge summaries for 300 randomly selected MIMIC-III notes  [14] and release scripts to process and use this data. To be fair to LLMs with different context lengths, we select MIMIC notes with no more than 4,000 tokens (measured with tiktoken’s p50_base tokenizer [^1] https://github.com/openai/tiktoken). As it is derived from MIMIC, this benchmark data should not be used outside of research contexts.
Models
We evaluate 6 different LLMs: GPT-4 [3] gpt4-0314, GPT-3.5 [15] gpt-3.5-turbo-0613, both accessed securely through the Azure OpenAI API. We also evaluate LLaMA-2 [7B,70B] [2], Mistral-7B [16], and Mixtral (8x7B) [17]. We evaluate each LLM both for generating summaries and for scoring summaries. We compare methods for scoring with three popular metrics: BERT-Score [18], and ROUGE-1/ROUGE-L [19].
Figure 2: Distribution of scores across different attributes. Attribute Structuring enables auditing which aspects of discharge summaries are generally accurate (e.g. Discharge test) and which are often inaccurate (e.g. Hospital consults).
Annotations and metrics
On a subset of our benchmark, three blinded human annotators evaluate the similarities between the extracted text for each attribute. They score 17 attributes for 30 total documents (20 test summaries generated by GPT-4 and 10 test summaries generated by Mixtral), for a total of 510 attribute labels. Each label is scored on a 4-point scale: Not similar, Somewhat similar, Very similar, Essentially the same. We evaluate the match between the manually annotated attribute labels and the automated attribute labels using Pearson correlation, Spearman (rank) correlation, and root mean-squared-error (RMSE).
3 Results
Main result: AS improves evaluation quality
Figure 3: Example interpretation of document similarity. Scores for each element of the ontology can be analyzed and grounded in quotes from the source summaries for better understanding.
Table 1: Main result: Attribute Structuring improves the match between automated scoring metrics and human annotators. For all methods, attribute structuring is performed by prompting GPT-4.
RMSE ↓ Pearson correlation Spearman correlation
Default/AS Default/AS Default/AS
ROUGE-1 0.17 / 0.17 0.36 / 0.64 0.34 / 0.48
ROUGE-L 0.26 / 0.18 0.35 / 0.59 0.32 / 0.44
BERT-Score 0.48 / 0.18 0.17 / 0.71 0.23 / 0.69
GPT-4 0.18 / 0.07 0.46 / 0.84 0.38 / 0.77
Table 2: Automatic evaluation scores using different LLMs for generating summaries and for evaluating. The Benchmark column uses Mixtral for structuring summaries as well as evaluating them, making it fully reproducible using open source tools. The remaining columns use GPT-4 for structuring and different LLMs for scoring the attributes. The relative rankings of different LLMs generators is preserved across different columns.
Benchmark Scorer (with GPT-4 structuring)
Mixtral (8x7B) AVG LLaMA-2 (7B) Mistral (7B) LLaMA-2 (70B) Mixtral (8x7B) GPT-3.5 GPT-4
Generator LLaMA-2 (7B) 36.2 45.97 65.4 47.1 49.0 38.5 40.3 35.5
Mistral 37.4 46.37 66.1 47.7 49.2 38.7 40.8 35.7
LLaMA-2 (70B) 38.4 49.70 66.7 50.6 54.3 43.4 43.6 39.6
Mixtral (8x7B) 39.2 51.30 68.4 52.4 54.9 46.0 44.9 41.2
GPT-3.5 39.4 52.78 73.1 53.7 56.7 45.9 45.7 41.6
GPT-4 45.1 56.02 73.8 56.9 61.2 51.2 48.8 44.2
Table 1 shows how well different scoring methods match the annotations provided by humans. Attribute Structuring yields a considerable improvement for all metrics (except for ROUGE-1 when measured with RMSE). Scoring with GPT-4 yields the best match with human annotators. The Pearson correlation coefficient of 0.84 for GPT-4 closely approaches the mean inter-annotator Pearson correlation of 0.90, and the Spearman correlation coefficient of 0.77 similarly nears the mean inter-annotator Spearman correlation of 0.87. The enhanced correlations across all metrics substantiate that attribute-based, fine-grained scoring outperforms holistic methodologies.
Fig. 2 shows how similarities break down across attributes for different attributes / metrics. Some attributes, such as Discharge test are more often correctly identified in the test summaries, whereas others, such as Hospital consults are rarely correct. Aggregating these attribute evaluations through AS allows for improved evaluation compared to evaluating an entire summary at once.
Summaries from stronger LLMs yield higher AS benchmark scores
We now evaluate how well different LLMs perform under an automatic evaluation using Attribute Structuring. The first column of Table 2 shows the benchmark scores achieved by different models; the benchmark scores rely only on the open-source Mixtral model to perform both structuring and scoring, enabling full reproducibility. GPT-4 achieves the highest score, followed by GPT-3.5 and Mixtral (8x7B).
We additionally evaluate how consistent the benchmark scores are with scores provided using different language models for structuring and scoring. The right section of Table 2 shows evaluation scores when structuring with GPT-4 and scoring with various different LLMs. We find that the absolute scores in different columns vary widely, but the relative ordering of LLM generators is almost perfectly preserved between the benchmark and all other settings (the only exception is that the Mixtral scorer slightly prefers the Mixtral generator to the GPT-3.5 generator).
Analysis: Attribute Structuring facilitates interpretation
Going beyond aggregate interpretations, Fig. 3 shows an example of how similarity can be audited for a single pair of summaries. Unlike posthoc feature importance [20, 21], or intrinsically interpretable models [22, 23], our interpretable grounding comes directly from an LLM, similar to recent works that use LLMs to generate explanations [24, 25] and ground those explanations in evidence [26, 27]. Since AS decomposes similarity into a series of attributes, the attributes can each be audited by directly prompting an LLM (here, GPT-4) to find a text span that serves as evidence for the extracted attribute in each document. Fig. 3 shows text spans for two discrepancies: Admission diagnosis and Pertinent lab tests, which are only partially retained in the test summary.
4 Discussion
Attribute Structuring constitutes an important step towards safely deploying LLMs in healthcare settings. As LLMs continue to generally improve, their ability to produce and evaluate clinical summaries with AS is likely to improve as well. Going forward, Attribute Structuring can be harnessed in a variety of ways to improve clinical NLP summarization, e.g. for evaluating free-form text such as conversations [28], improving model distillation by helping to filter effective summaries [29, 30], and improving the explainability of the process [31].
Limitations
One limitation of Attribute Structuring is that it incurs a high computational cost as multiple LLM calls are chained together; however, these costs may continue to decrease as models become more efficient [32]. Another limitation is that LLMs (and thereby Attribute Structuring) continue to be sensitive to prompts, increasing the need for methods to make LLMs more amenable to prompting [15, 33, 34] and to make finding strong prompts easier [35, 36, 37, 38]. Additionally, AS relies on a human-given ontology for evaluation, but could be made more general by developing methods to automatically define extract and define an ontology for a given domain. While Attribute Structuring helps to improve the automatic evaluation of summaries, there are nevertheless risks present with fully relying on LLMs without human supervision.
Appendix A Appendix
Here we include the prompts we used to:
  • Generate summaries
  • Structure to extract attributes ,and
  • Score each pair of attributes.
  • A prompt to score without structuring as a baseline.
1. Generation prompt:
Prompt = "You are tasked with generating a high quality clinical discharge summary for the provided input text INPUT_DOC. The summary has to be very relevant to the input document and cover the most important aspects of the input. Follow the following steps:
Step 1: Generate the most common sections that usually appear in a clinical discharge summary. Step 2: Not all the sections from Step 1 will have content. See which of those sections have contents from the INPUT_DOC. Remove the sections that don’t have information.
Step 3 : Use the following concepts to generate appropriate content. These are not Sections. Concepts: patient information and service type, diagnosis given at the time of admission, brief summary of initial presentation and diagnostic evaluation, pertinent physical findings relevant to diagnoses, goals of care; level of treatment,code status(e.g. curative,life-prolonging palliative, and symptomatic palliative), course in hospital; synotpic,problem-based description of sequential events and respective evaluations, treatments, and prognoses, hospital consults; description of specialty and/or allied health consults, procedures in hospital; a list of procedures with key findings and date, principal discharge diagnosis or main reason for admission and all additional pertinent diagnoses where applicable, discharge medications with specific description of new, altered, and discontinued medications and rationale for changes, lab tests and investigative results, tests ordered during the hospitalization that are pending at the time of discharge, outcome of care/condition at discharge; sense of the patient health status at discharge includes functional status, and cognitive status, outstanding issues for follow-up and recommendations to a recipient health-care provider during discharge, appointments after discharge including person responsible for scheduling, care provider, discharge instructions; list of information/education provided to the patient during discharge, main author of the discharge summary or attending clinician.
Step 4: Put the generated content in a coherent order. Format in such a way that the dishcarge summary has an excellent coherence, fluency, and consistency. Remember to use the sections you generated in Step 1, don’t make the concepts in Step 2 as a section, they are just suggestive concepts not sections.
Step 5: Remove sections with no information. Dont put ’not specified’ or ’not mentioned’ or ’none specified’ in a section. Just remove everything for that section including the section header.
Step 5: Return the final discharge summary with all the remaining sections that have contents. Remember to remove sections with no information. Context : "
2. Structuring prompt:
We use ResponseSchema from langchain to inform the model what each attribute description is.
ad_diag_schema
= ResponseSchema(name = ’ad_diag’, description=’preliminary or working diagnosis given at the time of admission’)
dc_diag_schema
= ResponseSchema(name = ’dc_diag’, description=’the list of principal discharge diagnosis or main reason for admission and all additional pertinent diagnoses where applicable’)
main_diag_schema
= ResponseSchema(name = ’main_diag’, description=’diagnosis mostly accountable for the largest portion of the patient"s stay, responsible for the greatest part of the length of stay ’)
history_schema
= ResponseSchema(name = ’history’, description=’a brief summary of initial presentation and diagnostic evaluation’)
physical_schema
= ResponseSchema(name = ’physical’, description=’pertinent physical findings relevant to diagnoses ’)
goals_schema
=ResponseSchema(name = ’goals’, description=’goals of care; level of treatment,code status(e.g. curative,life-prolonging palliative, and symptomatic palliative)’)
course_schema
= ResponseSchema(name = ’course’, description=’course in hospital; synotpic,problem-based description of sequential events and respective evaluations, treatments, and prognoses ’)
consults_schema
= ResponseSchema(name = ’consults’, description=’hospital consults; description of specialty and/or allied health consults’)
procedures_schema
= ResponseSchema(name = ’procedures’, description=’procedures in hospital; a list of procedures with key findings and date’)
ds_med_schema
= ResponseSchema(name = ’ds_med’, description=’a list of all discharge medications with specific description of new, altered, and discontinued medications and rationale for changes’)
lab_schema
= ResponseSchema(name = ’lab’, description=’pertinent lab tests and investigative results’)
ds_test_schema
= ResponseSchema(name = ’ds_test’, description=’tests ordered during the hospitalization that are pending at the time of discharge ’)
ds_status_schema
= ResponseSchema(name = ’ds_status’, description=’outcome of care/condition at discharge; sense of the patient health status at discharge includes functional status, and cognitive status’)
followup_schema
= ResponseSchema(name = ’followup’, description=’outstanding issues for follow-up and recommendations to a recipient health-care provider during discharge’)
appt_schema
= ResponseSchema(name = ’appt’, description=’appointments after discharge including person responsible for scheduling, care provider ’)
instruct_schema
= ResponseSchema(name = ’instruct’, description=’discharge instructions; list of information/education provided to the patient during discharge’)
author_schema
= ResponseSchema(name = ’author’, description=’main author of the discharge summary or attending clinician’)
prompt = "You are an expert in information extraction and structuring from clinical notes. Given a clinical note, create a structured output. For a given variable, if you can not determine/find a value, return NONE. Dont add any extra text, just the structured value. Here is the note: INPUT_DOC. format_instructions "
3. Structured attribute Scoring Prompt:
prompt = "You will be given a python dictionary containing a clinical variable as a key and a list containing two values for the variable as values.
Your task is to rate how similar the values are given the variable. Compare value1 and value2 for semantic similarity(similarity in meaning) given the variable and the criteria below. Two values can be very similar in meaning even if they are phrased differently. Also remember that this is a clinical document, take that into account. When scoring the similarity between two clinical terminologies, assign a value from 1 to 4, where 1 signifies a lack of similarity and 4 indicates identical meanings. Consider the context, clinical relevance, and semantic alignment between the terminologies. If the terminologies convey vastly different meanings, assign a score of 1. A score of 2 is appropriate for terminologies that, while related, represent different concepts or emphasize distinct elements. A score of 3 should be used for terminologies with substantial semantic overlaps but minor differences. Finally, if the terminologies are semantically equivalent and interchangeable, warranting no clinical distinction, assign a score of 4. Ensure that the assessment reflects the degrees of similarity in meaning, irrespective of syntactical differences, context, or minor variances in expression.
Please make sure you read and understand these instructions carefully. Return the similarity score. Return only the score, don’t include any other extra text. "
4. Document scoring without structuring prompt
prompt = """You will be given two clinical discharge summaries, summary1 and summary2.
Your task is to rate how similar the two summaries are from 1-10. 1 is not similar while 10 is essentially identical. Focus on the following important clinical variables when performing the comparison: How similar is the diagnosis, how similar are the goals, how similar is hospital course and history, how similar are the medications administered, how similar are the physical condition diagnoses, how similar are the followup consults and procedures, how similar are the lab tests performed, how similar are the patients discharge status, is the follow-up instructions similar, are there any similar appointments, and instructions only return the score, dont add any extra text Here are the inputs discharge summaries: summary1: [value1] summary2: [value2]. """
A.1 Prompts given to human annotators
Fig. A1 shows a sample interface given to human annotators. Human annotators were presented values for attributes through a web interface along with the prompt “For each attribute listed on the left, two possible values are provided(different colors). Please rate the semantic similarity between the values.”. Annotators are three machine learning researchers that are authors of the current study.
Figure A1: Annotation interface screenshot.